OpenStudy (anonymous):

(-3m)^2(2m^7)^2 Mulitply.

OpenStudy (saifoo.khan):

I have done this before as well. :O

OpenStudy (anonymous):

you have?

OpenStudy (saifoo.khan):

\[9m^2 \times 4m^{14}=36m^{2+14}=36m^{16}\]

OpenStudy (saifoo.khan):

Sorry that was a typo.

OpenStudy (anonymous):

hmmm. so what would it be?

OpenStudy (saifoo.khan):

nvm.

OpenStudy (saifoo.khan):

where are u stuck?? i can help u through!!

OpenStudy (saifoo.khan):

you can take my email and contact me: einstein.saifoo@gmail.com i do Mathlab as well.